# Harmonized Speculative Sampling

### v1

@@ 4.2 Effectiveness & Ablation Study
In this section, we first evaluate the effectiveness of HASS by comparing it with existing speculative sampling methods on speedup ratio and acceptance length. Then, we conduct ablation studies on harmonized objective distillation and context alignment.
@@ 4.2.2 Ablation Study on Harmonized Objective Distillation
We first study the effects of different K and the weight $w$ of the Top-K loss by varying these hyper-parameters and summarize the results in Figure 4. Training with the Top-K loss ($w>0$) always improves performance compared to training without the Top-K loss ($w=0$). HASS achieves the largest acceptance length when $w=0.5$. A small value of K ($K=1$) may result in performance degeneration, as the draft model only focuses on the token with the highest probability and consequently neglects other potential tokens. With a larger K, the Top-K loss generally brings better results, while the acceptance length is the largest when $K=5$.
We then conduct an experiment with LLaMA2-Chat 7B, where the fixed training dataset is replaced by the dataset generated by the target LLM. We observe that when using non-greedy decoding, the acceptance length increases from 4.92 to 5.19 averaging over three datasets. Therefore, information obtained from harmonized objective distillation is not equivalent to direct distillation from teacher-generated data. A better loss function may exist than Equation 1 to exploit the target LLM further. We leave this topic in future works.
@@ 4.2.3 Ablation Study on Harmonized Context Alignment
As the first training step of HASS is the same as EAGLE-2, we continually train EAGLE-2’s draft model weights with the Top-K loss and consider it the baseline. Without harmonized context alignment (EAGLE-2 + Top-K), the draft model performs the worst across all datasets. Training with 3/4 steps of harmonized context alignment generally obtains the most considerable acceptance length. When training with five steps of context alignment, the acceptance length decreases. We believe this is caused by the draft model’s limited capacity, as it predicts less accurately on former steps’ tokens when paying too much attention to the latter ones. Figure 5 shows the acceptance rate $\alpha$ across speculation steps on MT-bench following Li et al. (2024b). In later speculation steps, HASS performs better acceptance rates than EAGLE-2, demonstrating the effectiveness of harmonized context alignment.
Figure 5: Acceptance rates $\alpha$ of HASS and EAGLE-2 across different speculation steps on MT-bench with temperature T $\in\{0,1\}$.
@@ 5 Related Work
There have been a number of works on improving the acceptance rate of speculative sampling while maintaining the target distribution. Most of them fall into two categories. (1) The former category is aligned training that tries to obtain draft models aligned with the target LLM before the decoding stage. Zhou et al. (2023) propose a knowledge distillation approach and studies several strategies to improve the alignment. Li et al. (2024a) demonstrates that hidden states of the target LLM as input of the draft model provide extra feature uncertainty information. Xiao et al. (2024) also utilizes hidden states of the target LLM and introduces an RNN-based draft model design that achieves a comparable acceptance rate. GLIDE (Du et al., 2024) instead reuses kv-cache of the target LLM. It also notices the context misalignment when using information from the target LLM, but the proposed blockwise attention mask method can not solve the misalignment completely. (2) The latter category is efficient decoding, which designs sophisticated decoding strategies to utilize concurrency efficiently. Miao et al. (2024) proposes to utilize multiple draft models and designs a tree-based attention mechanism to verify multiple draft sequences efficiently. Li et al. (2024b) introduces a dynamic structure to save computation by pruning inefficient paths in the draft tree. Sun et al. (2024) studies improving the verification stage through optimal transportation. However, these works tend to only consider one of training and decoding, ignoring the linkage of these two stages. This work instead aims to link training and decoding, leading to harmonized speculative sampling.
@@ 6 Conclusion
This paper introduces HASS, a harmonized speculative sampling solution that addresses the misalignment between training and decoding on their objectives and contexts. Compared to its closest baseline, EAGLE-2, HASS improves the acceptance rate without any inference overhead. Experiments conducted on LLaMA2-Chat 7/13B, and LLaMA3-Instruct 8B demonstrate the effectiveness and efficiency of HASS. Averaging on MT-bench, HumanEval, and GSM-8K, HASS is 2.81x-3.65x faster than vanilla auto-regressive decoding, 8%-15% faster than EAGLE-2.

### v2

@@ 4.2 Effectiveness & Ablation Study
In this section, we first evaluate the effectiveness of HASS by comparing it with existing speculative sampling methods on acceptance length and speedup ratio. Then, we conduct ablation studies on harmonized objective distillation and harmonized context alignment.
@@ 4.2.2 Ablation Study on Harmonized Objective Distillation
We first study the effects of different K and the weight $w$ of the Top-K loss by varying these hyper-parameters and summarize the results in Figure 4. Training with the Top-K loss ($w>0$) always improves performance compared to training without the Top-K loss ($w=0$). HASS achieves the largest acceptance length when $w=0.5$. A small value of K may result in performance degeneration, as the draft model only focuses on the token with the highest probability and consequently neglects other potential tokens. With a larger K, the Top-K loss generally brings better results, while the acceptance length is the largest when K $=5$.
Loss Function Temperature = 0 Temperature = 1 Mean
Top-K Loss 4.99 4.84 4.92
Top-P Loss 5.03 4.76 4.90
Normed Top-K Loss (Linear) 4.97 4.83 4.90
Normed Top-K Loss (Softmax) 4.98 4.72 4.85
Bi-directional Top-K Loss 4.99 4.72 4.86
Recall@k Surrogate Loss 4.97 4.76 4.87
BiLD Loss 5.04 4.75 4.90
Table 3: Acceptance lengths $\tau$ of HASS with different kinds of loss functions for harmonized objective distillation. The results are conducted on LLaMA2-Chat 7B over the MT-bench dataset with temperature T $\in\{0,1\}$.
Since the harmonized objective distillation can be implemented with any loss function that focuses on the most probable tokens w.r.t. the target distribution, we further consider the following loss functions and compare them with the Top-K Loss:
  • Top-P Loss, where the $\hat{\Omega}$ is formed by the most probable tokens whose cumulative probability is just larger than P.
  • Normed Top-K Loss, where the target and draft distributions are both normalized over $\hat{\Omega}$. The normalization operation can be either linear or softmax.
  • Bi-directional Top-K Loss, where the distillation is conducted over the most probable tokens w.r.t. the target distribution as well as the draft distribution.
  • Recall@k Surrogate Loss (Patel et al., 2022), where a smooth approximation of the recall metric is obtained and is differentiable for direct optimization.
  • BiLD Loss (Li et al., 2024a), where the internal logits ranking information is captured by constructing logits differences with long-tail noise filtered out.
After searching the optimal hyper-parameters for each of the compared loss functions, we summarize their best results in Table 3. BiLD loss outperforms other loss functions under temperature T $=0$, while Top-K loss outperforms others under temperature T $=1$. Generally, Top-K loss shows the best performance. A better loss function may exist than Top-K loss to exploit the target LLM further. We leave this topic in future works.
We also conduct an experiment with LLaMA2-Chat 7B, where the fixed training dataset is replaced by the dataset generated by the target LLM (see Appendix 8). We observe that when using non-greedy decoding, the acceptance length increases from 4.92 to 5.19 averaging over three datasets. Therefore, information obtained from harmonized objective distillation is not equivalent to direct distillation from target-model-generated data.
@@ 4.2.3 Ablation Study on Harmonized Context Alignment
As the first training step of HASS is the same as EAGLE-2, we continually train EAGLE-2’s draft model weights with the Top-K loss and consider it the baseline. Without harmonized context alignment (EAGLE-2 + Top-K), the draft model performs the worst across all datasets. Training with 3/4 steps of harmonized context alignment generally obtains the most considerable acceptance length. When training with five steps of context alignment, the acceptance length decreases. We believe this is caused by the draft model’s limited capacity, as it predicts less accurately on former steps’ tokens when paying too much attention to the latter ones. Figure 5 shows the acceptance rate $\alpha$ across speculation steps on the MT-bench dataset following Li et al. (2024c). In later speculation steps, HASS performs better acceptance rates than EAGLE-2, demonstrating the effectiveness of harmonized context alignment.
Figure 5: Acceptance rates $\alpha$ of HASS and EAGLE-2 across different speculation steps on the MT-bench dataset with temperature T $\in\{0,1\}$.
As shown in Figure 5, the acceptance rates of HASS decrease compared to EAGLE-2 on LLaMA2-Chat 13B and LLaMA3-Instruct 70B at the first step (0-$\alpha$). The draft models degrade on the first speculation step with much attention paid to the latter speculation steps, while the first step’s acceptance rates are crucial to larger acceptance lengths. We conduct experiments to emphasize the significance of former speculation steps by reweighting the training loss from each step with a factor $\beta$. In specific, the step $j$’s training loss will be multiplied by $\beta^{j-1}$. Table 5 and Figure 6 show the acceptance lengths and acceptance rates of HASS with different reweight factors on LLaMA3-Instruct 70B over the MT-bench dataset, respectively. With the factor $\beta$ decreasing from 1.0 to 0.5, HASS achieves better acceptance lengths with different temperatures. Correspondingly, we perceive that the acceptance rate at the first speculation step is consistently higher with a smaller $\beta$, while the acceptance rates at the latter speculation steps generally decline. When the factor $\beta$ decreases to 0.3, too much emphasis is assigned to the first speculation step, leading to deterioration in acceptance length. Since the further exploration of an appropriate trade-off between different speculation steps is out of this paper’s scope, we leave it for future work.
Reweight Factor $\beta$ T = 0 T = 1 Mean
1.0 (Default) 4.62 4.59 4.61
0.7 4.65 4.61 4.63
0.5 4.67 4.62 4.65
0.3 4.65 4.61 4.63
Table 5: Acceptance lengths $\tau$ of HASS with different reweight factors $\beta$ for harmonized context alignment. The results are conducted on LLaMA3-Instruct 70B over the MT-bench dataset with temperature T $\in\{0,1\}$.
Figure 6: Acceptance rates $\alpha$ of HASS with different reweight factors $\beta$ for harmonized context alignment. The results are conducted on LLaMA3-Instruct 70B over the MT-bench dataset with temperature T $\in\{0,1\}$.
@@ 5 Related Work
There have been a number of works on improving the acceptance rate of speculative sampling while maintaining the target distribution. Most of them fall into two categories. (1) The former category is aligned training that tries to obtain draft models aligned with the target LLM before the decoding stage. Zhou et al. (2023) propose a knowledge distillation approach and study several strategies to improve the alignment. Li et al. (2024b) demonstrate that hidden states of the target LLM as input of the draft model provide extra feature uncertainty information. Xiao et al. (2024) also utilize hidden states of the target LLM and introduce an RNN-based draft model design that achieves a comparable acceptance rate. GLIDE (Du et al., 2024) instead reuses the KV cache of the target LLM. It also notices the context misalignment when using information from the target LLM, but the proposed blockwise attention mask method can not solve the misalignment completely. (2) The latter category is efficient decoding, which designs sophisticated decoding strategies to utilize concurrency efficiently. Miao et al. (2024) propose to utilize multiple draft models and design a tree-based attention mechanism to verify multiple draft sequences efficiently. Li et al. (2024c) introduce a dynamic structure to save computation by pruning inefficient paths in the draft tree. Sun et al. (2024) study improving the verification stage through optimal transportation. However, these works tend to only consider training or decoding, ignoring the linkage of these two stages. This work instead aims to link training and decoding, leading to harmonized speculative sampling.
@@ 6 Conclusion
This paper introduces HASS, a harmonized speculative sampling solution that addresses disharmonies between training and decoding on their objectives and contexts. Compared to its closest baseline, EAGLE-2, HASS improves the acceptance rate without any inference overhead. Experiments conducted on LLaMA2-Chat 7/13B and LLaMA3-Instruct 8/70B demonstrate the effectiveness and efficiency of HASS. Averaging on MT-bench, HumanEval, and GSM8K, HASS is 2.81x-4.05x faster than vanilla auto-regressive decoding, 8%-20% faster than EAGLE-2.
Appendix for ‘Learning Harmonized Representations for Speculative Sampling’
Appendix A Implementation of Harmonized Context Alignment
We present the pseudo code of harmonized context alignment, which is implemented without the customized attention mask, for better understanding. The actual implementation in our experiments is achieved by the customized attention mask as shown in Figure 3.
1def train_batch(
2 draft_model, # draft model
3 lm_head, # language model head
4 optimizer, # optimizer
5 forward_num, # aligning steps in harmonized context alignment
6 hidden_states_target, # target LLM’s feature
7 input_ids, # input tokens
8 ):
9 hidden_states_draft_list = []
10 for forward_idx in range(forward_num):
11 optimizer.zero_grad()
12 predict = draft_model(hidden_states_target, input_ids, hidden_states_draft_list)
13 hidden_states_draft = torch.cat([hidden_states_target[:, :1], predict[:, :-1]], dim=1).detach()
14 hidden_states_draft_list.append(hidden_states_draft)
15 target_head, pred_head = lm_head(hidden_states_target), lm_head(predict)
16 loss = feature_loss(hidden_states_target, predict) + logit_loss(target_head, pred_head)
17 loss.backward()
18 optimizer.step()
1def attention(
2 hidden_states_target, # target LLM’s feature
3 attention_mask, # causal attention mask
4 hidden_states_draft_list, # list of draft model’s features
5 ):
6 bs, seq_len = hidden_states_target.shape[0], hidden_states_target.shape[1]
7 query = q_proj(hidden_states_draft_list[-1]) if hidden_states_draft_list else q_proj(hidden_states_target)
8 key_t, value_t = k_proj(hidden_states_target), v_proj(hidden_states_target)
9 attn_weight = torch.matmul(query, key_t.transpose(2, 3)) / math.sqrt(query.shape[-1]) + attention_mask
10 indices = torch.arange(seq_len)
11 for i, hidden_states_draft in enumerate(hidden_states_draft_list[::-1]):
12 key_d, ind_q, ind_k = k_proj(hidden_states_draft), indices[i:], indices[:seq_len - i]
13 attn_weight_d = torch.matmul(query, key_d.transpose(2, 3)) / math.sqrt(query.shape[-1])
14 attn_weight[:, :, ind_q, ind_k] = attn_weight_d[:, :, ind_q, ind_k]
15 attn_weight_normed = F.softmax(attn_weight, dim=-1)
16 attn_output = torch.matmul(attn_weight_normed, value_t)
17 for i, hidden_states_draft in enumerate(hidden_states_draft_list[::-1]):
18 value_d, ind_q, ind_k = v_proj(hidden_states_draft), indices[i:], indices[:seq_len - i]
19 attn_output[:, :, ind_q] += attn_weight[:, :, ind_q, ind_k][..., None] * (value_d[:, :, ind_k] - value_t[:, :, ind_k])
20 attn_output = o_proj(attn_output.transpose(1, 2).reshape(bs, seq_len, -1))
21 return attn_output
Appendix B Harmonized Context Alignment on Tokens
In this section, we attempt to verify whether applying token alignment as well as feature alignment brings better performance. In specific, we use the tokens generated by the draft model for training in harmonized context alignment instead of using the tokens from training data. We apply feature and token alignment to EAGLE-2’s draft model weights and summarize the results in Table 6 and Figure 7.
Temperature = 0 Temperature = 1 Mean
EAGLE-2 4.44 4.23 4.34
Feature Only 4.83 4.60 4.72
Feature + Token (0.1) 4.81 4.57 4.69
Feature + Token (0.2) 4.78 4.51 4.65
Feature + Token (1.0) 4.28 4.11 4.20
Table 6: Acceptance lengths $\tau$ of applying feature and token alignment to EAGLE-2’s draft model weights, where ‘Token ($x$)’ denotes tokens from training data are replaced by draft-model-generated tokens with a probability of $x$. The results are conducted on LLaMA2-Chat 7B over the MT-bench dataset with temperature T $\in\{0,1\}$.
Figure 7: Acceptance rates $\alpha$ of applying feature and token alignment to EAGLE-2’s draft model weights, where ‘Token ($x$)’ denotes tokens from training data are replaced by draft-model-generated tokens with a probability of $x$. The results are conducted on LLaMA2-Chat 7B over the MT-bench dataset with temperature T $\in\{0,1\}$.
Feature only alignment brings the best performance, while token alignment leads to degeneration. With the probability of applying token alignment increasing from 0.1 to 1.0, the acceptance length decreases consistently. As shown in Figure 7, more token alignment generally causes lower acceptance rates. As a result, training with tokens generated by the draft model in harmonized context alignment hurts the acceleration performance.
Appendix C Hyper-Parameters of Top-K Loss
We conduct an ablation study on Top-K loss’s hyper-parameters, i.e., K and $w$, in Section 4.2.2 and show the averaged acceptance lengths over three datasets in Figure 4. Here, we present the speedup ratios and acceptance lengths of HASS with varied Ks and $w$s in Table 7.
MT-bench HumanEval GSM8K Mean
K $w$ Speedup $\tau$ Speedup $\tau$ Speedup $\tau$ Speedup $\tau$
T=0 1 1.0 2.89x 4.94 3.24x 5.19 3.11x 5.10 3.08x 5.08
5 1.0 2.90x 5.00 3.44x 5.29 3.33x 5.18 3.22x 5.16
10 1.0 2.99x 4.99 3.41x 5.29 3.32x 5.17 3.24x 5.15
50 1.0 2.85x 5.01 3.46x 5.29 3.41x 5.17 3.24x 5.16
100 1.0 2.93x 5.00 3.45x 5.29 3.45x 5.18 3.28x 5.16
10 0.0 2.77x 4.93 3.38x 5.22 3.18x 5.11 3.11x 5.09
10 0.1 2.98x 4.96 3.40x 5.26 3.51x 5.16 3.30x 5.13
10 0.2 2.87x 4.98 3.41x 5.29 3.35x 5.16 3.21x 5.14
10 0.5 3.00x 5.02 3.32x 5.31 3.50x 5.18 3.27x 5.17
10 2.0 2.94x 4.98 3.37x 5.29 3.34x 5.17 3.22x 5.15
T=1 1 1.0 2.58x 4.70 2.79x 4.80 2.83x 4.95 2.73x 4.82
5 1.0 2.64x 4.81 3.13x 4.94 2.93x 5.02 2.90x 4.92
10 1.0 2.70x 4.84 3.13x 4.91 2.87x 5.01 2.90x 4.92
50 1.0 2.62x 4.77 3.01x 4.88 2.99x 5.03 2.87x 4.89
100 1.0 2.66x 4.74 3.14x 4.97 2.90x 5.03 2.90x 4.91
10 0.0 2.61x 4.71 2.76x 4.84 2.79x 4.96 2.72x 4.84
10 0.1 2.69x 4.75 3.05x 4.94 2.87x 5.00 2.87x 4.90
10 0.2 2.66x 4.75 3.16x 4.95 2.88x 5.01 2.90x 4.90
10 0.5 2.68x 4.80 3.15x 4.93 2.96x 5.03 2.93x 4.92
10 2.0 2.68x 4.75 3.11x 4.89 2.89x 5.03 2.89x 4.89
Table 7: Speedup ratios and acceptance lengths $\tau$ of HASS with varied Ks and $w$s of the Top-K loss on LLaMA2-Chat 7B over MT-bench, HumanEval, and GSM8K datasets with temperature T $\in\{0,1\}$.
Appendix D Self-Distillation
Following existing speculative sampling methods (Zhou et al., 2023; Cai et al., 2024), we further use target-model-generated outputs to distill the draft model from the target model’s real output distribution, dubbed as self-distillation. In specific, we feed the prompts from the ShareGPT dataset into the target models recursively and collect the responses as multi-turn conversations for self-distillation. To study the effect of self-distillation, we conduct experiments by training the draft model with fixed data or model-generated data and summarize the results in Table 8.
MT-bench HumanEval GSM8K Mean
Model Data Speedup $\tau$ Speedup $\tau$ Speedup $\tau$ Speedup $\tau$
T=0 L2 7B F 2.99x 4.99 3.41x 5.29 3.32x 5.17 3.24x 5.15
MG 3.13x 5.25 3.85x 5.70 3.40x 5.57 3.46x 5.51
L2 13B F 3.23x 5.13 4.24x 6.05 3.48x 5.55 3.65x 5.58
MG 3.34x 5.27 4.42x 6.00 3.63x 5.61 3.80x 5.63
T=1 L2 7B F 2.70x 4.84 3.13x 4.91 2.87x 5.01 2.90x 4.92
MG 2.75x 4.97 3.39x 5.24 3.13x 5.35 3.09x 5.19
L2 13B F 3.28x 4.98 3.78x 5.86 3.37x 5.41 3.48x 5.42
MG 3.33x 5.02 3.76x 5.80 3.60x 5.42 3.56x 5.41
Table 8: Speedup ratios and acceptance lengths of HASS with fixed or target-model-generated training data. F and MG stand for ‘Fixed’ and ‘Model-Generated’, respectively. L2 represents LLaMA2-Chat.
On LLaMA2-Chat 7B, self-distillation consistently brings improvements. On LLaMA2-Chat 13B, self-distillation only achieves marginally better or comparable results, as the corresponding draft model’s performance is saturated. Especially, the acceptance lengths of the self-distilled draft model are lower than that of the vanilla one on the HumanEval dataset. It may be due to the code generation dataset HumanEval being less similar to the training dataset ShareGPT compared with MT-bench and GSM8K.
Appendix E Drafting Hyper-Parameters
Li et al. (2024c) find that the draft token’s confidence score is strongly positively correlated with the acceptance rate, and accordingly propose the context-aligned dynamic draft tree, which can be dynamically adjusted with two hyper-parameters: ‘depth’ and ‘number of tokens’. ‘Depth’ decides the draft tree’s depth during the expansion phase, while ‘number of tokens’ decides how many draft tokens will be kept during the reranking phase. Increasing both these hyper-parameters surely leads to a larger acceptance length. Nevertheless, sending more draft tokens into the target model for verification causes a higher overhead in real applications. Therefore, we vary these hyper-parameters and report the speedup ratios in Table 9 to find a better trade-off.
Depth 5 6 7 8 9
# Tokens 40 60 80 100 40 60 80 100 40 60 80 100 40 60 80 100 40 60 80 100
T=0 L2 7B EAGLE-2 2.48x 2.78x 2.61x 2.69x 2.69x 2.66x 2.70x 2.79x 2.71x 2.86x 2.91x 2.95x 2.60x 2.60x 2.88x 2.89x 2.28x 2.54x 2.55x 2.65x
HASS-MG 3.09x 3.02x 3.04x 3.22x 3.08x 3.13x 3.19x 3.22x 3.14x 3.11x 3.29x 3.27x 3.07x 3.16x 3.31x 3.32x 2.78x 2.75x 2.95x 3.03x
L2 13B EAGLE-2 2.63x 2.96x 3.04x 3.06x 3.01x 3.02x 3.18x 3.22x 2.78x 3.12x 3.14x 3.24x 2.98x 3.12x 3.19x 3.26x 2.49x 2.64x 2.69x 2.72x
HASS-MG 3.25x 3.31x 3.24x 3.25x 3.33x 3.34x 3.49x 3.40x 3.19x 3.42x 3.36x 3.40x 3.15x 3.40x 3.40x 3.37x 2.70x 2.74x 3.09x 3.02x
T=1 L2 7B EAGLE-2 2.31x 2.37x 2.55x 2.36x 2.42x 2.39x 2.33x 2.40x 2.49x 2.66x 2.65x 2.44x 2.36x 2.48x 2.38x 2.64x 2.29x 2.22x 2.27x 2.42x
HASS-MG 2.79x 2.89x 2.86x 2.88x 2.72x 2.75x 2.92x 2.82x 2.83x 2.76x 2.81x 2.75x 2.49x 2.68x 2.77x 2.77x 2.30x 2.35x 2.58x 2.50x
L2 13B EAGLE-2 2.92x 3.11x 2.88x 2.79x 3.06x 3.04x 3.16x 2.93x 3.05x 3.14x 3.14x 3.11x 3.00x 3.13x 3.15x 2.98x 2.61x 2.72x 2.65x 2.54x
HASS-MG 3.24x 3.30x 3.27x 3.19x 3.33x 3.33x 3.40x 3.28x 3.19x 3.26x 3.24x 3.26x 3.15x 3.26x 3.19x 3.17x 2.62x 2.77x 2.74x 2.84x
Table 9: Speedup ratios of EAGLE-2 and HASS-MG with varied depths and numbers of tokens on the MT-bench dataset with temperature T $\in\{0,1\}$, where HASS-MG denotes HASS trained with self-distillation. L2 represents LLaMA2-Chat.
[…]